# A Context for Manifold Calculus

### v2

@@ 4. Classification of Homogeneous Functors
Let $k\geq 0$, let $M$ be a manifold, let $\mathcal{C}$ be a pointed $\infty$-category with small colimits, and let $F:N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{C}$ be a functor which is polynomial of degree $\leq k$. We say that $F$ is homogeneous of degree $k$ if $F$ evaluates to a zero object at each element $U\in\operatorname{Disj}\left(M\right)$ with fewer than $k$ components, or equivalently, if the $\left(k-1\right)$th level $T_{k-1}F$ of a Taylor cotower of $F$ (Example 3.19) carries every object to a zero object. For example, every functor of polynomial $\leq 0$ is homogeneous of degree 00; a $1$-excisive isotopy functor is homogeneous of degree $1$ if and only if it is a cosheaf (by Theorem 3.10 and Example 3.8).
We are interested in homogeneous functors for the following reason: Suppose that $F$ is an isotopy functor. A cofiber of $T_{k-1}F\to T_{k}F$, called a $k$th homogeneous layer of $F$, is then an isotopy functor which is homogeneous of degree $\leq k$: It is an isotopy functor by Lemma 3.5; it is polynomial of degree $\leq k$ by Theorem 3.10; homogeneity is clear. Therefore, if $F$ has a convergent Taylor cotower, then up to extension problem we can reduce the analysis of $F$ to those of the homogeneous layers of $F$.
In this section, we prove that homogeneous functors are classified by unordered configuration spaces and give an interpretation of the homogeneous layer in terms of higher derivatives.
Notation 4.1.
Let $M$ be a manifold, and let $\mathcal{C}$ be a pointed $\infty$-category with small colimits. We let $\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)$ denote the full subcategory of $\operatorname{Fun}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ spanned by the exhaustive isotopy functors that are homogeneous of degree $k$.
Recall that if $M$ is an $n$-manifold, then $I_{M}^{=k}$ denotes the (non-full) subcategory of $\operatorname{Disj}\left(M\right)$ spanned by the isotopy equivalences between elements of $\operatorname{Disj}\left(M\right)$ with exactly $k$ components (Notation 2.30). The nerve of $I_{M}^{=k}$ classifies homogeneous functors in the following sense:
@@
$\theta:\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)\xrightarrow{\simeq}\operatorname{Fun}_{\mathrm{istp}}\left(N\left(I_{M}^{=k}\right),\mathcal{C}\right).$
@@
$p:\operatorname{Exc}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)\to\operatorname{Fun}_{\mathrm{istp}}\left(N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right),\mathcal{C}\right)$
is a trivial fibration. Let $\operatorname{Fun}^{\prime}_{\mathrm{istp}}\left(N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right),\mathcal{C}\right)\subset\operatorname{Fun}_{\mathrm{istp}}\left(N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right),\mathcal{C}\right)$ denote the full subcategory spanned by the functors which carry every object in $N\left(\operatorname{Disj}^{\leq k-1}\left(M\right)\right)$ to a zero object. The functor
$\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)\to\operatorname{Fun}^{\prime}_{\mathrm{istp}}\left(N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right),\mathcal{C}\right)$
is a pullback of $p$, so it is a trivial fibration. On the other hand, Proposition 2.34 shows that the functor
$\operatorname{Fun}^{\prime}_{\mathrm{istp}}\left(N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right),\mathcal{C}\right)\to\operatorname{Fun}_{\mathrm{istp}}\left(N\left(I_{M}^{=k}\right),\mathcal{C}\right)$
@@
Now if $M$ is a manifold and $k\geq 0$, then a functor on $N\left(I_{M}^{=k}\right)$ is an isotopy functor if and only if it maps every morphism of $N\left(I_{M}^{=k}\right)$ to an equivalence. Therefore, Proposition 4.2 says that homogeneous functors of degree $k$ on a manifold $M$ can be classified in terms of the classifying space of $I_{M}^{=k}$, i.e., a localization of $N\left(I_{M}^{=k}\right)$ with respect to all morphisms [Lur24, Tag 01MY]. The following proposition identifies the homotopy type of this classifying space.
@@
$N\left(I_{M}^{=k}\right)\xrightarrow{\simeq}\operatorname{Sing}B_{k}\left(M\right)$
@@
The identification of the homotopy type of $N\left(I_{M}^{=k}\right)$ dates back at least to [Wei99, Lemma 3.5] and has been reproved again and again (e.g., [AF15, Lemma 2.12]). The point of Proposition 4.3 is that there is an explicit weak homotopy equivalence, whose account seem to be lacking in the existing literature.
@@
The proposition is a consequence of the following assertion:
  • ($\ast$)
    Let $P\subset I_{M}^{=1}$ be a full subcategory which forms a basis of the topology on $M$. Then there is a weak homotopy equivalence
    $N\left(P\right)\to\operatorname{Sing}M$
    which carries each object $U\in N\left(P\right)$ to a point in $U$.
To see how ($\ast$) implies the proposition, notice that each element $U\in I_{M}^{=k}$ determines an element $B^{\prime}_{k}\left(U\right)\in I_{B_{k}\left(M\right)}^{=1}$, where $B^{\prime}_{k}\left(U\right)$ denotes the subset of $B_{k}\left(M\right)$ consisting of the points $\{p_{1},\dots,p_{k}\}\in B_{k}\left(M\right)$ which intersects every component of $U$. (This is consistent with Notation 2.30.) The map $U\mapsto B^{\prime}_{k}\left(U\right)$ identifies $I_{M}^{=k}$ with a full subcategory of $I_{B_{k}\left(M\right)}^{=1}$ which form a basis of the topology on $B_{k}\left(M\right)$. We can then apply ($\ast$) to $P=I_{M}^{=k}$ to obtain the proposition.
To prove ($\ast$), let $n$ denote the dimension of $M$. Let $\operatorname{Chart}\left(P,M\right)$ denote the category whose elements are the embeddings $\iota:\mathbb{R}^{n}\to M$ such that $\iota\left(\mathbb{R}^{n}\right)\in P$, and whose morphisms $\iota\to\iota^{\prime}$ are the continuous maps rendering the diagram
[…]
commutative. The forgetful functor $N\left(\operatorname{Chart}\left(P,M\right)\right)\to N\left(P\right)$ is a trivial fibration. It will therefore suffice to construct a weak homotopy equivalence $N\left(\operatorname{Chart}\left(P,M\right)\right)\to\operatorname{Sing}M$ which carries each object $\iota\in N\left(\operatorname{Chart}\left(P,M\right)\right)$ to the point $\iota\left(0\right)\in\operatorname{Sing}M$.
Given integers $p,q\geq 0$ and continuous maps $\sigma:\left|\Delta^{p}\right|\to\mathbb{R}^{n}$ and $\tau:\left|\Delta^{q}\right|\to\mathbb{R}^{n}$, we define a map $\sigma\star\tau:\left|\Delta^{p+q+1}\right|\to\mathbb{R}^{n}$ by
$\sigma\left(s_{0},\dots,s_{p},t_{0},\dots,t_{q}\right)=\begin{cases}s\sigma\left(\frac{s_{0}}{s},\dots,\frac{s_{p}}{s}\right)+t\tau\left(\frac{t_{0}}{t},\dots,\frac{t_{q}}{t}\right)&\text{if }s,t\neq 0,\\ \tau\left(t_{0},\dots,t_{q}\right)&\text{if }s=0,\\ \sigma\left(s_{0},\dots,s_{p}\right)&\text{if }t=0,\end{cases}$
where we set $s=\sum_{i=0}^{p}s_{i}$ and $t=\sum_{i=1}^{q}t_{i}$. Note that if $f:\Delta^{p^{\prime}}\to\Delta^{p}$ and $g:\Delta^{q^{\prime}}\to\Delta^{q}$ are morphisms of simplicial sets, then $\left(\sigma\star\tau\right)\left(\left|f\star g\right|\right)=\left(\sigma\left|f\right|\right)\star\left(\tau\left|g\right|\right)$.
Define
$\Phi_{M}:N\left(\operatorname{Chart}\left(P,M\right)\right)\to\operatorname{Sing}M$
as follows: Let $\sigma=\left(\iota_{0}\xrightarrow{f_{1}}\cdots\xrightarrow{f_{p}}\iota_{p}\right)$ be a $p$-simplex of $N\left(\operatorname{Chart}\left(P,M\right)\right)$. Define $\alpha_{\sigma}:\left|\Delta^{p}\right|\to\mathbb{R}^{n}$ inductively by
$\alpha_{\sigma}=\begin{cases}0&\text{if }p=0,\\ \left(f_{p}\alpha_{d_{p}\sigma}\right)\star 0&\text{if }p>0.\end{cases}$
One can show by induction on $p$ that for each $p>0$ and $0\leq i\leq p$, we have
$\alpha_{\sigma}\partial_{i}=\begin{cases}f_{p}\alpha_{d_{p}\sigma}&\text{if }i=p,\\ \alpha_{d_{i}\sigma}&\text{if }i<p,\end{cases}$
and that for each $0\leq i\leq p$ we have $\alpha_{\sigma}\sigma_{i}=\alpha_{s_{i}\sigma}$. Therefore, the assignment $\sigma\mapsto\iota_{p}\circ\alpha_{\sigma}$ determines a morphism of simplicial sets $N\left(\operatorname{Chart}\left(P,M\right)\right)\to\operatorname{Sing}M$, which we denote by $\Phi_{M}$. We claim that $\Phi_{M}$ is a weak homotopy equivalence.
For each $U\in P$, let $P_{U}\subset P$ denote the full subcategory spanned by the elements that are contained in $U$. Consider the diagram
$SSF:P\to\SS$
given by $U\mapsto N\left(\operatorname{Chart}\left(P_{U},U\right)\right)$. The colimit of this diagram is $N\left(\operatorname{Chart}\left(P,M\right)\right)$, so there is a commutative diagram
${{\operatorname{hocolim}_{U\in P}N(\operatorname{Chart}(P_{U},U))}}$${{N(\operatorname{Chart}(P,M))}}$${{\operatorname{hocolim}_{U\in P}\operatorname{Sing}(U)}}$${{\operatorname{Sing}(M).}}$$\scriptstyle{h}$$\scriptstyle{\Phi_{M}}$$\scriptstyle{g}$$\scriptstyle{f}$
Here $f$ is induced by the maps $\{\Phi_{U}\}_{U\in P}$. We must show that the map $\Phi$ is a weak homotopy equivalence. Since each $N\left(\operatorname{Chart}\left(P_{U},U\right)\right)$ has a terminal object, the map $f$ is a weak homotopy equivalence. The map $h$ is also a weak homotopy equivalence by [Lur17, Theorem A.3.1]. It will therefore suffice to show that the map $g$ is a weak homotopy equivalence. For this, it suffices to show that the diagram $F$ is projectively cofibrant. We will show that, in fact, for each pair of simplicial subsets $A\subset B\subset N\left(\operatorname{Chart}\left(P,M\right)\right)$, the map $A\cap F\to B\cap F$ is a projective cofibration. (Here $A\cap F$ denotes the diagram $SSP\to\SS$ given by $U\mapsto A\cap F\left(U\right)$.) Working simplex by simplex, it will suffice to consider the case where there is a pushout square
[…]
Depict $\sigma$ as $\iota_{0}\xrightarrow{f_{1}}\cdots\xrightarrow{f_{p}}\iota_{p}$, where each $\iota_{i}$ is an embedding $\mathbb{R}^{n}\to M$. Let $SSj:P^{\mathrm{op}}\to\operatorname{Fun}\left(P,\SS\right)$ denote the Yoneda embedding. Then the square
${{\partial\Delta^{p}\times j(\iota_{p}(\mathbb{R}^{n}))}}$${{A\cap F}}$${{\Delta^{p}\times j(\iota_{p}(\mathbb{R}^{n}))}}$${{B\cap F}}$
is cocartesian. The left vertical arrow is a projective cofibration, so the right vertical arrow is also a projective cofibration. The proof is now complete. ∎
Combining Propositions 4.2 and 4.3, we obtain:
@@
$\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)\xrightarrow{\simeq}\operatorname{Fun}_{\mathrm{istp}}\left(N\left(I_{M}^{=k}\right),\mathcal{C}\right)\xleftarrow{\simeq}\operatorname{Fun}\left(\operatorname{Sing}B_{k}\left(M\right),\mathcal{C}\right).$
Propositions 4.2 and 4.3 lead to an interesting observation if we apply it to homogeneous layers of isotopy functors [^4] The author learned this interpretation from [Mun10].. Let $n\geq 0$, let $k\geq 1$, let $M$ be an $n$-manifold, let $\mathcal{C}$ be a pointed $\infty$-category with small colimits, and let $F:N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{C}$ be an isotopy functor. Let $L_{k}F$ denote a cofiber of the natural transformation $T_{k-1}F\to T_{k}F$ and let $\overline{L}_{k}F:\operatorname{Sing}B_{k}\left(M\right)\to\mathcal{C}$ denote the corresponding functor. Let us compute the value of $\overline{L}_{k}F$ at a point$\{p_{1},\dots,p_{k}\}\in B_{k}M$. First, choose pairwise disjoint open sets $B_{1},\dots,B_{k}\subset M$ that are homeomorphic to $\mathbb{R}^{n}$, and set $U=B_{1}\cup\dots\cup B_{k}$. Then

### v3

@@ 3. Classification of Homogeneous Functors
In this section, we classify homogeneous functors, the building blocks of Taylor towers in manifold calculus. More formally, let $k\geq 0$, let $M$ be a manifold, let $\mathcal{C}$ be a pointed $\infty$-category with small colimits, and let $F:\operatorname{Open}\left(M\right)\to\mathcal{C}$ be a functor which is polynomial of degree $\leq k$. We say that $F$ is homogeneous of degree $k$ if $F$ evaluates to a zero object at each element $U\in\operatorname{Disj}\left(M\right)$ with fewer than $k$ components; or equivalently, if the $\left(k-1\right)$th level $T_{k-1}F$ of a Taylor cotower of $F$ (Example 2.2) carries every object to a zero object. For example, every functor of polynomial of degree $\leq 0$ is homogeneous of degree 0; a $1$-excisive exhaustive isotopy functor is homogeneous of degree $1$ if and only if it is a cosheaf (by Theorem 1.9 and Example 1.7).
We are interested in homogeneous functors for the following reason: Suppose that $F$ is an isotopy functor. A cofiber of $T_{k-1}F\to T_{k}F$, called a $k$th homogeneous layer of $F$, is then an isotopy functor which is homogeneous of degree $\leq k$: It is an isotopy functor by Lemma 1.4; it is polynomial of degree $\leq k$ by Theorem 1.9; homogeneity is clear. Therefore, if $F$ has a convergent Taylor cotower, then we can reduce the analysis of $F$ to those of the homogeneous layers of $F$, at least up to extension problems.
In this section, we prove that homogeneous functors are classified by unordered configuration spaces and “objects of sections trivial near the fat diagonal” (Corollary 3.5 and Theorem 3.16). To illustrate the utility of the classification, we also give a connectivity estimate of these approximations (Corollary 3.17).
Notation 3.1.
Let $M$ be a manifold, and let $\mathcal{C}$ be a pointed $\infty$-category with small colimits. We let $\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)$ denote the full subcategory of $\operatorname{Fun}\left(\operatorname{Open}\left(M\right),\mathcal{C}\right)$ spanned by the exhaustive isotopy functors that are homogeneous of degree $k$.
Recall that if $M$ is an $n$-manifold, then $I_{M}^{=k}$ denotes the (non-full) subcategory of $\operatorname{Disj}\left(M\right)$ spanned by the isotopy equivalences between the elements of $\operatorname{Disj}\left(M\right)$ with exactly $k$ components (Notation A.7). The poset $I_{M}^{=k}$ classifies homogeneous functors in the following sense:
@@
$\theta:\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)\xrightarrow{\simeq}\operatorname{Fun}_{\mathrm{istp}}\left(I_{M}^{=k},\mathcal{C}\right).$
@@
$p:\operatorname{Exc}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)\to\operatorname{Fun}_{\mathrm{istp}}\left(\operatorname{Disj}^{\leq k}\left(M\right),\mathcal{C}\right)$
is a trivial fibration. Let $\operatorname{Fun}^{\prime}_{\mathrm{istp}}\left(\operatorname{Disj}^{\leq k}\left(M\right),\mathcal{C}\right)\subset\operatorname{Fun}_{\mathrm{istp}}\left(\operatorname{Disj}^{\leq k}\left(M\right),\mathcal{C}\right)$ denote the full subcategory spanned by the functors which carry every object in $\operatorname{Disj}^{\leq k-1}\left(M\right)$ to a zero object. The functor
$\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)\to\operatorname{Fun}^{\prime}_{\mathrm{istp}}\left(\operatorname{Disj}^{\leq k}\left(M\right),\mathcal{C}\right)$
is a pullback of $p$, so it is a trivial fibration. On the other hand, Proposition B.19 shows that the functor
$\operatorname{Fun}^{\prime}_{\mathrm{istp}}\left(\operatorname{Disj}^{\leq k}\left(M\right),\mathcal{C}\right)\to\operatorname{Fun}_{\mathrm{istp}}\left(I_{M}^{=k},\mathcal{C}\right)$
@@
Now if $M$ is a manifold and $k\geq 0$, then a functor on $I_{M}^{=k}$ is an isotopy functor if and only if it maps every morphism of $I_{M}^{=k}$ to an equivalence. Therefore, Proposition 3.2 says that homogeneous functors of degree $k$ on a manifold $M$ can be classified in terms of the classifying space of $I_{M}^{=k}$, i.e., a localization of $I_{M}^{=k}$ with respect to all morphisms [Lur25, Tag 01MY]. The following proposition identifies the homotopy type of this classifying space.
@@
$I_{M}^{=k}\xrightarrow{\simeq}\operatorname{Sing}B_{k}\left(M\right)$
@@
The identification of the homotopy type of $I_{M}^{=k}$ dates back at least to [Wei99, Lemma 3.5] and has been reproved again and again (e.g., [AF15, Lemma 2.12]). We include a proof for later reference.
@@
We consider the maps
$I_{M}^{=k}\xleftarrow{\pi}\int^{U\in I_{M}^{=k}}\operatorname{Sing}B^{\prime}_{k}\left(U\right)\xrightarrow{\phi}\operatorname{Sing}B_{k}\left(M\right),$
where $B^{\prime}_{k}\left(-\right)$ is defined in Notation A.7. The map $\pi$ is a trivial fibration because it is a left fibration with contractible fibers. So it has a section $\sigma$, and $\phi\circ\sigma$ carries each element $U\in I_{M}^{=k}$ to a point in $B^{\prime}_{k}\left(U\right)$. It will therefore suffice to show that $\phi$ is a weak homotopy equivalence. According to (the argument of) [Lur09, Corolary 3.3.4.6], it will suffice to show that the map
$\operatorname{hocolim}_{U\in I_{M}^{=k}}\operatorname{Sing}B^{\prime}_{k}\left(U\right)\to\operatorname{Sing}B_{k}\left(M\right)$
is a weak homotopy equivalence. This follows from [Lur17, Theorem A.3.1]. ∎
Combining Propositions 3.2 and 3.3, we obtain:
@@
$\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)\xrightarrow{\simeq}\operatorname{Fun}_{\mathrm{istp}}\left(I_{M}^{=k},\mathcal{C}\right)\xleftarrow{\simeq}\operatorname{Fun}\left(\operatorname{Sing}B_{k}\left(M\right),\mathcal{C}\right).$
Remark 3.6.
[…]
@@
In the situation of Corollary 3.5, it is relatively easy to pass from $\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)$ to $\operatorname{Fun}\left(\operatorname{Sing}B_{k}\left(M\right),\mathcal{C}\right)$: We just have to restrict a homogenous functor to $I_{M}^{=k}$, and then pass to the localization. We now explain how to go in the reverse direction (Theorem 3.16).
The starting point is the observation that for a manifold $M$, the equivalence $\operatorname{Fun}\left(\operatorname{Sing}B_{k}\left(M\right),\mathcal{C}\right)\xrightarrow{\simeq}\operatorname{Fun}_{\mathrm{istp}}\left(I_{M}^{=k},\mathcal{C}\right)$ coming from Proposition 3.3 is given by
$F\mapsto\left(U\mapsto\operatorname{colim}_{\operatorname{Sing}B^{\prime}_{k}\left(U\right)}F\right).$
(This follows from the proof of Proposition 3.3. Also, we used the informal notation of Remark B.6, applied to the cocartesian fibration $\int^{U\in I_{M}^{=k}}\operatorname{Sing}B^{\prime}_{k}\left(U\right)$). A bit more consideration then leads to the following proposition:
Proposition 3.7.
Let $M$ be a manifold, let $\mathcal{C}$ be a pointed $\infty$-category with small colimits, let $k\geq 0$, and let $F:\operatorname{Sing}B_{k}\left(M\right)\to\mathcal{C}$ be a functor. Define another functor $\widetilde{F}:\operatorname{Open}\left(M\right)\to\mathcal{C}$ by
$\widetilde{F}\left(U\right)=\operatorname{colim}_{\operatorname{Sing}B_{k}\left(U\right)}F.$
Then:
  1. (1)
    $\widetilde{F}$ is $k$-excisive, exhaustive, and isotopy invariant.
  2. (2)
    The assignment $F\mapsto L_{k}\widetilde{F}$ determines an inverse equivalence
    $\operatorname{Fun}\left(\operatorname{Sing}B_{k}\left(M\right),\mathcal{C}\right)\xrightarrow{\simeq}\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)$
    of the equivalence of Proposition 3.2.
We need a few lemmas for the proof of Proposition 3.7.
Notation 3.8.
Let $M$ be a manifold and let $k\geq 0$. We write $B^{\prime\prime}_{k}\left(M\right)$ for the complement of $B^{\prime}_{k}\left(M\right)$ in $B_{k}\left(M\right)$. That is, $B^{\prime\prime}_{k}\left(M\right)$ is the subspace of $B_{k}\left(M\right)$ consisting of the points $\{p_{1},\dots,p_{k}\}$ such that the map $\{p_{1},\dots,p_{k}\}\to\pi_{0}\left(M\right)$ is not injective.
Lemma 3.9.
Let $U$ be a manifold whose components are homeomorphic to $\mathbb{R}^{n}$, and let $k\geq 1$. The map
$\int^{V\in\operatorname{Disj}^{\leq k-1}\left(U\right)}\operatorname{Sing}B_{k}\left(V\right)\to\operatorname{Sing}\left(B^{\prime\prime}_{k}\left(U\right)\right)$
is a weak homotopy equivalence.
Proof.
By [Lur09, Lemma 3.3.4.1 and Proposition 3.3.4.2], it suffices to show that the diagram $\{\operatorname{Sing}B_{k}\left(V\right)\to\operatorname{Sing}B^{\prime\prime}_{k}\left(U\right)\}_{V\in\operatorname{Disj}^{\leq k-1}\left(U\right)}$ is a homotopy colimit diagram of simplicial sets. According to [Lur09, Theorem A.3.1], it suffices to show that, for each point $S\in B^{\prime\prime}_{k}\left(U\right)$, the full subcategory $\operatorname{Disj}^{\leq k-1}\left(U\right)_{S}\subset\operatorname{Disj}^{\leq k-1}\left(U\right)$ spanned by the objects $V\in\operatorname{Disj}^{\leq k-1}\left(U\right)$ such that $S\subset V$ is weakly contractible. Since $S$ lies in no more than $\left(k-1\right)$ components of $U$, such a $V$ necessarily lies in the union of the components of $U$ containing points in $S$. Thus $\operatorname{Disj}^{\leq k-1}\left(U\right)_{S}$ has a final object, and is in particular weakly contractible. ∎
Proof of Proposition 3.7.
The claim is trivial if $k=0$, so we will assume that $k\geq 1$ throughout.
We start with (1). By Proposition B.8, it suffices to show that the functor
$\operatorname{Sing}B_{k}\left(-\right):\operatorname{Open}\left(M\right)\to\mathcal{S}$
is $k$-excisive, exhaustive, and isotopy invariant. The first two follows from [Lur17, Theorem A.3.1], while isotopy invariance is obvious. [^4] Incidentally, $B_{k}\left(-\right)$ is one of the “classic” polynomial functors that appeared in the original paper of Weiss; see [Wei99, Example 2.4].
For (2), let $U\in I_{M}^{=k}$ and $F\in\operatorname{Fun}\left(\operatorname{Sing}B_{k}\left(M\right)\right)$. Since $B_{k}\left(U\right)$ is the disjoint union of $B^{\prime}_{k}\left(U\right)$ and $B^{\prime\prime}_{k}\left(U\right)$, we have an equivalence
$\left(\operatorname{colim}_{\operatorname{Sing}B^{\prime}_{k}\left(U\right)}F\right)\vee\left(\operatorname{colim}_{\operatorname{Sing}B^{\prime\prime}_{k}\left(U\right)}F\right)\xrightarrow{\simeq}F\left(U\right)$
natural in $F$ and $U$. (But beware that the summands on the right-hand side are not functors of $U\in\operatorname{Open}\left(M\right)$, because an inclusion of open sets of $M$ may fail to be injective on $\pi_{0}$.) Moreover, Lemma 3.9 give us equivalences
$T_{k-1}\widetilde{F}\left(U\right)$ $=\operatorname{colim}_{V\in\operatorname{Disj}^{\leq k-1}\left(U\right)}\operatorname{colim}_{\operatorname{Sing}B_{k}\left(V\right)}F$
$\simeq\operatorname{colim}_{\int^{V\in\operatorname{Disj}^{\leq k-1}\left(U\right)}\operatorname{Sing}B_{k}\left(V\right)}F\left(W\right)$
$\simeq\operatorname{colim}_{\operatorname{Sing}\left(B^{\prime\prime}_{k}\left(U\right)\right)}F.$
Under these equivalences, the map $T_{k-1}\widetilde{F}\left(U\right)\to\widetilde{F}\left(U\right)$ can be identified with the inclusion of the summand
$\left(\operatorname{colim}_{\operatorname{Sing}B^{\prime\prime}_{k}\left(U\right)}F\right)\to\left(\operatorname{colim}_{\operatorname{Sing}B^{\prime}_{k}\left(U\right)}F\right)\vee\left(\operatorname{colim}_{\operatorname{Sing}B^{\prime\prime}_{k}\left(U\right)}F\right).$
Hence we obtain equivalences
$L_{k}\widetilde{F}\left(U\right)\simeq\mathrm{cofib}\left(T_{k-1}\widetilde{F}\left(U\right)\to\widetilde{F}\left(U\right)\right)\simeq\operatorname{colim}_{\operatorname{Sing}B^{\prime}_{k}\left(U\right)}F$
natural in $F$ and $U$. As we observed right before stating Proposition 3.7, the right-hand side computes the desired inverse, so we are done. ∎
Proposition 3.7 has a refinement when there is a good supply of open sets of $B_{k}\left(M\right)$ of configurations where two or more points are “close.” To explain this in more detail, we need a few preliminaries.
Warning 3.10.
Following Warning 1.1, we will focus on manifold calculus of presheaves for the remainder of this section. If $M$ is a manifold and $\mathcal{C}$ is a pointed $\infty$-category with small limits, we write $\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M^{\mathrm{op}};\mathcal{C}\right)\subset\operatorname{Fun}\left(\operatorname{Open}\left(M\right)^{\mathrm{op}},\mathcal{C}\right)$ for the full subcategory spanned by the functors that are isotopy invariant, exhaustive, and homogeneous of degree $k$ when viewed as a functor $\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathcal{C}$.
Notation 3.11.
Let $\mathcal{C}$ be an $\infty$-category with small limits, let $X$ be a topological space, and let $F:\operatorname{Sing}X\to\mathcal{C}$ be a functor. For each subset $A\subset X$, we define $\Gamma\left(F;A\right)=\operatorname{lim}_{\operatorname{Sing}A}F.$